### v1

@@ 4.2 BETR Algorithm Workflow
Prior distribution. We impose zero-centered priors on the parameters, encoding the default assumption of no grade bias.
$\alpha\sim\mathcal{N}(0,\sigma_{a}^{2}),\quad\delta_{i}\sim\mathcal{N}^{+}(0,\tau^{2}),\ i\in\{B,C,D,E\}$ (5)
@@
Likelihood function. Given parameters $(\alpha,\delta)$, the likelihood of the observed pairwise preference data ($d^{+}\succ d^{-}$) is
$P(\mathcal{D}\mid\alpha,\delta)=\prod_{q\in\mathcal{Q}_{\mathrm{train}}}\prod_{(d^{+},d^{-})\in\mathcal{P}_{q}}\sigma\!\big(a\,\Delta s+u_{t^{+}}-u_{t^{-}}\big),$ (6)
@@
Posterior and MAP estimations. According to Bayes’ theorem[56], the posterior distribution is proportional to the product of the likelihood and priors:
$P(\alpha,\delta\mid\mathcal{D})\propto P(\mathcal{D}\mid\alpha,\delta)\cdot P(\alpha)\cdot P(\delta).$ (7)
@@
$\max_{\alpha,\ \delta_{B},\delta_{C},\delta_{D},\delta_{E}\geq 0}$ $\frac{1}{|\mathcal{Q}_{\mathrm{train}}|}\sum_{q\in\mathcal{Q}_{\mathrm{train}}}\frac{1}{\max(1,|\mathcal{P}_{q}|)}\sum_{(d^{+},d^{-})\in\mathcal{P}_{q}}\log\sigma\!\big(a\,\Delta s+u_{t^{+}}-u_{t^{-}}\big)$ (8)
$-\frac{1}{2\tau^{2}}\big(\delta_{B}^{2}+\delta_{C}^{2}+\delta_{D}^{2}+\delta_{E}^{2}\big)-\frac{1}{2\sigma_{a}^{2}}\alpha^{2}.$
@@
Training details. We split benchmark queries into non-overlapping sets: $\mathcal{Q}_{\mathrm{train}}$ for fitting calibrator parameters, and $\mathcal{Q}_{\mathrm{val}}$ for hyperparameter selection and early stopping via grid search. Splitting was query-grouped, ensuring that candidate windows for the same query did not span sets, thereby preventing information leakage. During BETR training, candidate sets $\mathcal{C}_{q}$ were constructed using the same candidate generation process $\mathrm{Cand}(\cdot)$ as at inference, reducing distribution shift.
Final reranking score. At inference time, for any candidate window $d\in\mathcal{C}_{q}$, the final BETR ranking score is
$r(q,d)=\hat{a}\,s(q,d)+\hat{u}_{\mathrm{Grade}(d)}.$ (9)
That is, the scale-calibrated semantic relevance score is combined with the corresponding evidence-grade bias.
4.3 PICO-extended Schema and Knowledge Graph Construction
Schema Definition and Constrained Extraction. Following Youtu-GraphRAG, the schema is defined as
$\mathcal{S}\triangleq\left\langle\mathcal{S}_{e},\mathcal{S}_{r},\mathcal{S}_{\text{attr}}\right\rangle,$ (10)
where $\mathcal{S}_{e}$, $\mathcal{S}_{r}$, and $\mathcal{S}_{\text{attr}}$ represent the sets of entity, relation, and attribute types, respectively. Based on this schema, the LLM extraction agent performs constrained triple extraction for each document (denoted by $x$):
$\mathcal{T}(x)=\Big\{(h,r,t),(e,r_{\text{attr}},e_{\text{attr}})\ \Big|\ \{f(h),f(t),f(e)\}\in\mathcal{S}_{e},\ \{r,r_{\text{attr}}\}\in\mathcal{S}_{r},\ e_{\text{attr}}\in\mathcal{S}_{\text{attr}}\Big\}.$ (11)
Here, $(h,r,t)$ denotes an entity–relation–entity triple, and $(e,r_{\text{attr}},e_{\text{attr}})$ an entity–attribute pair; $f(\cdot)$ maps text spans to types, constraining the extraction space and suppressing noise.
Instantiation of the PICO-extended schema. To explicitly incorporate the PICO framework into extraction constraints, we instantiate $\mathcal{S}$ as a collection of PICO-related types: (1) entity types $\mathcal{S}_{e}^{\text{PICO}}$: Population, Condition, Intervention, Comparator, Outcome, Timepoint, plus domain-specific extensions frequently occurring in sports rehabilitation literature (e.g., Arm, Device); (2) relation types $\mathcal{S}_{r}^{\text{PICO}}$: directed relations linking studies to PICO entities (e.g., has_population, has_condition, uses_intervention, compares_to, reports_outcome); (3) attribute types $\mathcal{S}_{\text{attr}}^{\text{PICO}}$: key dimensions refining PICO elements (e.g., age_bin, followup_weeks, measure_name, protocol_params). A complete list is provided in Supplementary Table S1.
@@ 4.3 PICO-extended Schema and Knowledge Graph Construction
Schema self-expansion. Youtu-GraphRAG allows agents to propose schema extensions based on document content:
@@
$\Delta\mathcal{S}=\left\langle\Delta\mathcal{S}_{e},\Delta\mathcal{S}_{r},\Delta\mathcal{S}_{\text{attr}}\right\rangle=\mathbb{I}\!\left[f_{\mathrm{LLM}}(x,\mathcal{S})\odot\mathcal{S}\right]\geq\mu,$ (12)
@@
where $\mu$ is the confidence threshold for accepting new schema elements, and $\Delta\mathcal{S}$ contains candidate extensions for entities, relations, and attributes, respectively. For our task, this mechanism enables graph and index updates across disease subtypes.
@@
Dually-perceived community compression and knowledge tree indexing. To reduce density and noise in the raw triple graph and shorten retrieval context, we adopted Youtu-GraphRAG’s dually-perceived community detection. The affinity between node $e_{i}$ and community $\mathcal{C}_{m}$ is defined as
@@
$\phi(e_{i},\mathcal{C}_{m})=\underbrace{\mathbb{S}_{r}(e_{i},\mathcal{C}_{m})}_{\text{relational}}\ \oplus\ \lambda\,\underbrace{\mathbb{S}_{s}(e_{i},\mathcal{C}_{m})}_{\text{semantic}},$ (13)
@@
where $\mathbb{S}_{r}$ measures Jaccard similarity over relation-type sets $\Psi(\cdot)$, $\mathbb{S}_{s}$ measures subgraph semantic similarity, and $\lambda$ is a trade-off coefficient. Iteratively, the algorithm performs pairwise community merging based on the following threshold criterion:
$\mathbb{E}\!\left[\phi\!\left(e_{i},\mathcal{C}_{a}^{(t)}\right)\right]-\mathbb{E}\!\left[\phi\!\left(e_{i},\mathcal{C}_{b}^{(t)}\right)\right]<\epsilon.$ (14)
@@
Ultimately, the community structure forms a four-layer knowledge tree $\mathcal{K}=\bigcup_{\ell=1}^{4}L_{\ell}$: $L_{4}$ = community, $L_{3}$ = keyword, $L_{2}$ = entity–relation triple, $L_{1}$ = attribute. This knowledge tree serves as a structured index for subsequent graph retrieval and community filtering.
4.4 SR-RAG Pipeline
As previously mentioned, SR-RAG introduced three key improvements to Youtu-GraphRAG: (1) a PICO-guided HyDE channel fused with graph retrieval via RRF; (2) two-stage reranking using ColBERT and a cross-encoder; (3) evidence-grade-aware retrieval that recalls guideline evidence separately to prevent dilution, followed by BETR-based reranking and a soft quota strategy to exclude semantically irrelevant candidates. For specific hyperparameter settings, see Supplementary Table S3.
PICO-guided HyDE retrieval channel. The original HyDE method addresses poor zero-shot dense retrieval performance without relevance annotations[27]. LLMs generate several hypothetical documents related to a user query, which are subsequently used for dense retrieval in the corpus. This paradigm tolerates factual errors in hypothetical documents: because retrieval relies on dense embeddings rather than lexical matching, semantic similarity is preserved even when generated content contains inaccuracies[27]. Subsequent work identified another advantage of HyDE: bridging the semantic gap between queries and documents[38]. In RAG, document-form inputs generally outperform raw question-form queries; thus, HyDE effectively improves retrieval quality and downstream performance. Based on this finding, we incorporated PICO soft constraints into the HyDE prompt: available P/I/C/O/T keywords were extracted from the query as anchors (missing fields are permitted). Hypothetical documents must reuse these anchors and are prohibited from fabricating missing fields. This design ensures HyDE serves purely as a retrieval intermediary for semantic alignment, without deviating from the original query’s PICO elements. The specific prompts are detailed in Supplementary Listing S1.
Two-stage reranking. We first used ColBERT as the coarse ranking model. Specifically, we used mxbai-edge-colbert-v0, which was recently proposed by Clavié et al.[57]. ColBERT uses the MaxSim mechanism for scoring:
Here, $\mathbf{e}_{i}^{q}$ and $\mathbf{e}_{j}^{d}$ denote the $i$-th and $j$-th token embeddings of the query and window, respectively. The MaxSim mechanism computes, for each query token, its maximum similarity with any window token and sums across tokens[29], improving sensitivity to medical abbreviations, scale names, and synonyms while increasing matching robustness[58].
After coarse ranking, we used BGE-reranker-v2-m3 as a cross-encoder to perform fine-grained ranking on the top-$K$ high-scoring candidates[59]. The cross-encoder jointly encodes the query and each candidate window via concatenation, enabling full interaction through multiple self-attention layers and outputting an overall relevance logit $s(q,d)$. The computational cost of the cross-encoder increases linearly with the number of candidates. By adopting a two-stage pipeline, the system reduces reranking time and computational cost while balancing precise term matching and semantic understanding of long texts.
Evidence-grade-aware retrieval and scoring. To ensure retrieval and ranking align with EBM principles, we explicitly introduced evidence-grade information at both the candidate generation and scoring stages. First, the corpus was partitioned by annotated evidence grade into Grade A and Grades B–E, forming two candidate pools. Candidate generation and truncation were performed separately on each pool. The candidate sets were then merged, and initial reranking yielded cross-encoder logit scores. BETR calibration was then applied for final global ranking:
$r(q,d)=\hat{a}\,s(q,d)+\hat{u}_{\mathrm{Grade}(d)}.$ (16)
This design enables the system to prioritize higher-grade evidence when candidate relevance is comparable.
4.5 Benchmark Construction
To facilitate automated evaluation of SR-RAG, we created 1,637 QA pairs using a combination of LLM generation and human review. The benchmark composition by sub-condition and guideline set is summarized in Supplementary Table S4. Given the unique logic of SR-RAG’s QA process and to ensure both reliable evaluation and cross-system reusability, we adopted the following pipeline.
First, using the corpus with completed evidence-grade annotation and chunking, we performed stratified sampling by evidence grade to obtain candidate evidence windows, filtering those with insufficient information density or lacking substantive clinical conclusions. For each candidate window, an LLM generated a clinical question strictly corresponding to the window’s core conclusion while simultaneously extracting PICOT elements. During generation, questions had to be directly supported by the window; inclusion of information not present in the window was prohibited[60]. Additionally, the LLM annotated evidence certainty (sufficient or uncertain) for each question, enabling stratified evaluation and error analysis.
Next, we performed retrieval accessibility checks for each generated question, similar to the round-trip consistency filtering in ARES[60]. Using a system-agnostic hybrid retrieval baseline (combining sparse and dense methods), we retrieved the top-$K$ candidate windows. If the seed window was recalled within top-$K$, the question was deemed “accessible” and entered the main split; otherwise, it entered the challenge split for manual review.
For reference evidence construction, we selected gold windows for each accessible question from a larger candidate pool. This entailed reranking candidates by relevance, then having the LLM grade the evidence relationship (strongly supportive/supportive/weakly related/unrelated). Strongly supportive and supportive windows formed the gold evidence set. Based on gold windows, we generated hierarchical reference answers: an exact answer summarizing the core conclusion, and an ideal answer organized by evidence grades (A–E), with Grade A prioritized and remaining evidence presented in descending pyramid order. All statements were strictly aligned with gold windows, with source identifiers for traceability. The LLM then decomposed the exact answer into atomic facts (nuggets) for automated assessment[30].
Finally, four sports science graduate students manually reviewed all QA pairs, with particular attention to samples marked “evidence uncertain” and those in the challenge split. Approximately 2,000 questions were generated initially; after human screening, 1,637 were retained as the official evaluation set. We extracted questions, exact answers, and nuggets and compiled them into a public dataset for reusability. Because some gold windows originate from institutionally subscribed materials, the public version excluded evidence window text.
4.6 Evaluation Process
Figure 5: Automated and human evaluation pipelines of SR-RAG.
SR-RAG evaluation comprised two dimensions: automated and manual (Fig. 5). For automated evaluation, we used all 1,637 QA pairs to conduct multi-metric assessments across generative models and ablation experiments. Metrics included: (1) Nugget coverage: extent to which the answer covers GT core factual units; (2) Answer faithfulness: whether answer statements are supported by retrieved evidence; (3) Semantic similarity: semantic consistency between answer and GT; (4) PICOT match accuracy: alignment of the answer with P/I/C/O/T elements.
Nugget coverage and PICOT match accuracy were implemented via LLM-as-judge[32]: for each nugget, the LLM determined whether it was covered by the system answer; for each extracted P/I/C/O/T field, the LLM evaluated whether it matched the reference (paraphrasing permitted; missing GT fields default to matched). Coverage and match rates were then aggregated. Answer faithfulness and semantic similarity were assessed via the RAGAS framework[20]. RAGAS first extracted claims from the answer, then determined for each claim whether it was supported by the actually retrieved context, yielding a faithfulness score (also LLM-as-judge based). Semantic similarity was computed as cosine similarity between text embeddings of the system answer and GT.
For manual evaluation, we invited five sports rehabilitation experts to review a random sample of 20 questions, each using identical review materials. Review materials include task instructions, scoring criteria, evaluation dimensions, and the question set. For each question, materials provide the question and GT, system-generated answer, retrieved evidence fragments, and a scoring sheet. A five-point Likert scale was used, with each question scored independently across five dimensions: (1) Medical factual accuracy: whether conclusions, values, and prescriptions align with current evidence, guidelines, and consensus; (2) Answer faithfulness: whether key statements are substantiated by retrieved fragments; (3) Answer relevance: whether the response addresses core concerns and covers key constraints; (4) Safety: whether the answer contains potentially harmful or misleading content; (5) PICOT alignment: whether the answer is structured around the question’s P/I/C/O/T elements without population or intervention mismatch. We aggregated scores from all five experts across five dimensions for the 20 questions and computed means and standard deviations per question and dimension for analysis.
4.7 Implementation details
All experiments were implemented in Python 3.11. Graph construction followed Youtu-GraphRAG’s schema-bounded extraction and community compression using NetworkX, with graph artifacts stored as JSON. Dense retrieval was implemented with sentence-transformers and FAISS; candidate lists from dense, graph, and HyDE retrieval were fused via RRF. Two-stage reranking used ColBERT(mxbai-edge-colbert-v0) for coarse ranking and a cross-encoder (bge-reranker-v2-m3) for final logits. BETR training used PyTorch 2.6.0 with the Adam optimizer (learning rate: 0.05, epochs: 80, $K=20$ negatives per positive). Hyperparameters ($\tau$) were selected via grid search on a query-disjoint split (train: nn=1,310; validation: nn=327). Unless otherwise specified, LLM decoding used temperature = 0.
Data availability
The SR-RAG benchmark dataset is publicly available at https://huggingface.co/datasets/Ning311/sr-rag-benchmark. The SR-RAG knowledge graph is publicly available at https://huggingface.co/datasets/Ning311/sr-rag-knowledge-graph. Because some gold windows originate from institutionally subscribed materials, the public release does not include the full text of evidence windows.
Code availability
The SR-RAG code will be publicly released after publication.
Acknowledgements
We thank the expert clinicians for participating in the human evaluation and providing valuable feedback. This work was supported by the National Key Research and Development Program of China (Grant No. 2022YFC3600201), the National Natural Science Foundation of China (Grant No. 32000838), and the Chinese Universities Scientific Fund (Grant No. 2024JNPD002).
Author contributions
J.Z., J.S., and Y.L. contributed to the conception and design of the study. T.S. and J.Z. designed the BETR algorithm. W.T., Z.L., J.Z., J.S., and Z.W. contributed to data collection and benchmark construction. J.X.L., J.L., and X.L. provided medical expertise and were responsible for the recruitment and execution of the expert clinician evaluation. Y.L. supervised the overall project, reviewed the experimental design and results, and acquired funding. J.Z. and Y.L. drafted and revised the manuscript and approved the final version.
Competing interests
The authors declare no competing interests.
Additional information
This preprint includes supplementary material.
Correspondence and requests for materials should be addressed to the corresponding author.
Supplementary Materials
S1. Schema Specifications
@@ Acknowledgements
Table S1 lists the PICO-extended schema used for graph construction and retrieval. Table S2 lists the PICO-neutral schema used in the w/o PICO-extended schema ablation.
@@
Table S1: PICO-extended schema used in graph construction and retrieval.
@@
Table S2: PICO-neutral schema used in the w/o PICO-extended schema ablation.

### v2

@@ B.3 Ordered Hierarchical Parameterization
B.4 Prior Distribution
We impose zero-centered priors on the parameters, encoding the default assumption of no grade bias:
$\alpha\sim\mathcal{N}(0,\sigma_{a}^{2}),\quad\delta_{i}\sim\mathcal{N}^{+}(0,\tau^{2}),$
$i\in\{B,C,D,E\}$ (B.3)
@@ B.4 Prior Distribution
B.5 Likelihood Function
Given parameters $(\alpha,\delta)$, the likelihood of the observed pairwise preference data ($d^{+}\succ d^{-}$) is
$P(\mathcal{D}\mid\alpha,\boldsymbol{\delta})=\prod_{q\in\mathcal{Q}_{\mathrm{train}}}\;\prod_{(d^{+},d^{-})\in\mathcal{P}_{q}}\\ \sigma\!\bigl(a\,\Delta s+u_{t^{+}}-u_{t^{-}}\bigr),$ (B.4)
@@ B.5 Likelihood Function
B.6 Posterior and MAP Estimation
According to Bayes’ theorem (Gelman et al., 2013), the posterior distribution is proportional to the product of the likelihood and priors:
$P(\alpha,\delta\mid\mathcal{D})\propto P(\mathcal{D}\mid\alpha,\delta)\cdot P(\alpha)\cdot P(\delta).$ (B.5)
@@ B.6 Posterior and MAP Estimation
$\begin{split}\max_{\alpha,\;\boldsymbol{\delta}\geq\mathbf{0}}\;&\frac{1}{|\mathcal{Q}_{\mathrm{tr}}|}\!\sum_{q\in\mathcal{Q}_{\mathrm{tr}}}\frac{1}{\max(1,|\mathcal{P}_{q}|)}\\ &\times\!\sum_{(d^{+},d^{-})\in\mathcal{P}_{q}}\!\log\sigma\!\bigl(a\,\Delta s+u_{t^{+}}\!-\!u_{t^{-}}\bigr)\\ &-\frac{1}{2\tau^{2}}\bigl(\delta_{B}^{2}\!+\!\delta_{C}^{2}\!+\!\delta_{D}^{2}\!+\!\delta_{E}^{2}\bigr)-\frac{\alpha^{2}}{2\sigma_{a}^{2}}.\end{split}$ (B.6)
@@ B.8 Training Details
We split the 1,637 benchmark queries into three non-overlapping sets: $\mathcal{Q}_{\mathrm{train}}$ ($n=983$) for fitting calibrator parameters, $\mathcal{Q}_{\mathrm{val}}$ ($n=327$) for hyperparameter selection and early stopping via grid search, and $\mathcal{Q}_{\mathrm{test}}$ ($n=327$) reserved exclusively for end-to-end evaluation. Splitting was query-grouped, ensuring that candidate windows for the same query did not span sets, thereby preventing information leakage. During BETR training, candidate sets $\mathcal{C}_{q}$ were constructed using the same candidate generation process $\mathrm{Cand}(\cdot)$ as at inference, reducing distribution shift.
B.9 BETR Training Curves and Learned Parameters
Figure B.1: BETR training curves showing MAP objective convergence and learned grade bias parameters.
Fig. B.1 shows the training dynamics of BETR. The learned parameters are summarized in Table B.1.
Table B.1: Learned BETR parameters.
Parameter Symbol Value
Scale $a=\exp(\alpha)$ 1.0348
Grade A bias $u_{A}$ 0 (anchor)
Grade B bias $u_{B}$ 0.1287-0.1287
Grade C bias $u_{C}$ 0.2575-0.2575
Grade D bias $u_{D}$ 0.3863-0.3863
Grade E bias $u_{E}$ 0.5151-0.5151
Shrinkage $\tau$ selected 1.0
Scale prior $\sigma_{a}$ fixed 5.0
The near-unity scale parameter ($a\approx 1.035$) confirms that BETR preserves the original semantic relevance scale. The grade bias decrements ($\approx 0.129$ per tier) form an approximately uniform staircase, consistent with the pyramid-shaped evidence hierarchy assumed in EBM.
Appendix C Schema Specifications and Community Compression
C.1 Schema Self-expansion
@@
Youtu-GraphRAG allows agents to propose schema extensions based on document content:
@@
$\Delta\mathcal{S}=\bigl\langle\Delta\mathcal{S}_{e},\;\Delta\mathcal{S}_{r},\;\Delta\mathcal{S}_{\text{attr}}\bigr\rangle$
$=\mathbb{I}\!\bigl[f_{\mathrm{LLM}}(x,\mathcal{S})\odot\mathcal{S}\bigr]\geq\mu,$ (C.1)
@@
where $\mu$ is the confidence threshold for accepting new schema elements, and $\Delta\mathcal{S}$ contains candidate extensions for entities, relations, and attributes. For our task, this mechanism enables graph and index updates across disease subtypes.
C.2 Dually-perceived Community Compression and Knowledge Tree Indexing
@@
To reduce density and noise in the raw triple graph and shorten retrieval context, we adopted Youtu-GraphRAG’s dually-perceived community detection. The affinity between node $e_{i}$ and community $\mathcal{C}_{m}$ is defined as
@@
$\phi(e_{i},\mathcal{C}_{m})=\underbrace{\mathbb{S}_{r}(e_{i},\mathcal{C}_{m})}_{\text{relational}}\ \oplus\ \lambda\,\underbrace{\mathbb{S}_{s}(e_{i},\mathcal{C}_{m})}_{\text{semantic}},$ (C.2)
@@
where $\mathbb{S}_{r}$ measures Jaccard similarity over relation-type sets $\Psi(\cdot)$, $\mathbb{S}_{s}$ measures subgraph semantic similarity, and $\lambda$ is a trade-off coefficient. Iteratively, the algorithm performs pairwise community merging based on the threshold criterion:
$\mathbb{E}\!\left[\phi\!\left(e_{i},\mathcal{C}_{a}^{(t)}\right)\right]-\mathbb{E}\!\left[\phi\!\left(e_{i},\mathcal{C}_{b}^{(t)}\right)\right]<\epsilon.$ (C.3)
@@
Ultimately, the community structure forms a four-layer knowledge tree $\mathcal{K}=\bigcup_{\ell=1}^{4}L_{\ell}$: $L_{4}$ = community, $L_{3}$ = keyword, $L_{2}$ = entity–relation triple, $L_{1}$ = attribute.
C.3 Schema Tables
@@
Table C.1 lists the PICO-extended schema used for graph construction and retrieval. Table C.2 lists the PICO-neutral schema used in the w/o PICO-extended schema ablation.
@@
Table C.1: PICO-extended schema used in graph construction and retrieval.
@@
Table C.2: PICO-neutral schema used in the w/o PICO-extended schema ablation.